WHAT IS THE SLOPE OF THE LINE THROUGH THE POINTS A(-1,6) AND B(5,2)?
Found 2 solutions by stanbon, jai_kos:
Answer by stanbon(48568)

WHAT IS THE SLOPE OF THE LINE THROUGH THE POINTS A(-1,6) AND B(5,2)?
----
slope of a line = [change in y values]/[change if corresponding x values}
In your case:
slope = [6-2]/[-1-5]=4/(-6) = -2/3
Cheers,
Stan H.

Answer by jai_kos(139)

Given two points A(-1,6) AND B(5,2)
Take it as x1 = -1, y1 = 6, x2 = 5, y2 = 2
The slope of the line given by m = (y2 -y1) / (x2 -x1)

m = (2 - 6) / (5 - (-1))
m = - 4 / 5 + 1
m = -4 / 6
m = -2 / 3
Therefore the slpoe is equal to (-2/3)